x, y, z are respectively the sines and p, q, r are respectively cosines of the angle α,β,¥ which are in AP with common difference 2π/3......then (1)yz+zx+xy is equal to....... (2) value of x^2(qy-rz) y^2(rz-px) +z^2(px-qy) is equal to...

x, y, z  are respectively the sines and p, q, r are respectively cosines of the angle α,β,¥ which are in AP with common difference 2π/3......then  (1)yz+zx+xy  is equal to....... (2) value of x^2(qy-rz) y^2(rz-px) +z^2(px-qy)  is equal to... 

Grade:11

1 Answers

Aditya Gupta
5 years ago
Let alpha= x-2pi/3 beta=x and Gama=x+2pi/3
xy+yz+zx can be found by writing alpha beta and gamma in terms of x and then applying the usual formula of compound angle. So xy+yz+zx= -3/4